a certain recipe requires 4 1/5 cups of flour and 2 3/7 cups of sugar. If 4/9 of the recipe is to be made, how much sugar is needed?
Answer by Boreal(15235)

Want 4/9 of 2 and 3/7
7/7=1; 14/7 =2; 17/7 = 2 and 3/7
(4/9)*(17/7)=(68/63) cups, just over 1 cup, and that cannot be reduced.